# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ Appendix H Impact of the sampler on the Gaussian likelihood posteriors
Figure 29: Comparison of posteriors obtained from a standard Gaussian likelihood analysis when using different MCMC samplers for the same noiseless mock data vector. In pink and in grey, we show the posteriors sampled with nautilus (Lange 2023) from a likelihood defined by a sample covariance based on mocks from the anisotropic systematics model and the standard isotropic systematics, respectively. In green and orange, we show posteriors sampled from the same two likelihoods as before, but sampling with the multinest sampler (Feroz et al. 2009) instead. The dashed black lines indicate the true cosmology of the input mock data vector generated from the anisotropic systematics model (see Table 4).
@@ Appendix I KiDS-SBI posteriors and parameter constraints
To expand on the results of the KiDS-SBI analysis of the KiDS-1000 gold sample presented in Fig. 19 and Table 3, we present the full posterior learnt in KiDS-SBI in Fig. 30 for both the anisotropic systematics and standard isotropic systematics models. Table 5 also provides the parameter best estimates for all cosmological and astrophysical parameters varied in the forward models. As shown in Fig. 30, both models leave $A_{\mathrm{bary}}$, $n_{\mathrm{s}}$ and $h_{0}$ unconstrained, while constraining $A_{\mathrm{IA}}$. The posteriors of both models strongly prefer values of $A_{\mathrm{IA}}$ greater than zero and less than 1.5. Despite the wide range of the $A_{\mathrm{IA}}$ priors, the posterior stays within a physical regime for the parameter which indicates a low chance of residual systematics strongly biasing the parameter constraints.
Figure 30: Posterior contours of the main constrained cosmological parameters from the KiDS-SBI analysis of the KiDS-1000 cosmic shear data assuming the anisotropic systematics model (in blue) compared against posterior contours from other analyses. In orange, we show the posterior from the same data while assuming the standard isotropic systematics model. In black and pink, we show the posteriors from the equivalent analyses done assuming a Gaussian likelihood for a given model. For clarity purposes, the $\bm{\delta}_{z}$ parameters are marginalised out, while $\omega_{\mathrm{b}}$ and $\omega_{\mathrm{c}}$ are folded into $\Omega_{\mathrm{m}}$. The green contour shows the posterior from the cosmic microwave background constraints from the TT,TE,EE+lowE modes (Planck Collaboration et al. 2020). Note that the Planck TT,TE,EE+lowE contours do not have any marginals in $A_{\mathrm{IA}}$ or $A_{\mathrm{bary}}$ as the CMB is not sensitive to the IAs of galaxies or baryonic feedback.

### v2

@@ Appendix G Impact of the sampler on the Gaussian likelihood posteriors
Figure 27: Comparison of posteriors obtained from a standard Gaussian likelihood analysis when using different MCMC samplers for the same noiseless mock data vector. In pink and in grey, we show the posteriors sampled with nautilus (Lange 2023) from a likelihood defined by a sample covariance based on mocks from the anisotropic systematics model and the standard isotropic systematics, respectively. In green and orange, we show posteriors sampled from the same two likelihoods as before, but sampling with the multinest sampler (Feroz et al. 2009) instead. All inference shown is done using the real KiDS data vector.
Appendix H Posterior variations with noise realisations
As discussed in Sect. 6, different uncertainty realisations of the anisotropic systematics model in KiDS-SBI at the same cosmology produces a significant random scatter in the measured final pseudo-Cls. Figure 28 shows how the choice of noise realisation impacts the inferred posteriors at a given cosmology. We find that overall the variations in the posteriors between uncertainty realisations are small with the mean marginal of $S_{8}$ fluctuating by up to ±$\pm 0.5\sigma$ and the $A_{\mathrm{IA}}$ marginal can fluctuate up to ±$\pm 0.75\sigma$.
Figure 28: Posterior contours from five different mock data vectors given the anisotropic systematics model within KiDS-SBI over the prior space shown in Table 18. Each data realisation assumes the same input parameters (as given by the black line and shown in Table 21) while varying the random seeds of the underlying matter fields, galaxy positions, intrinsic galaxy shapes, and shear biases. The posteriors shown in Fig. 25,14, 15, and 16 are based on the 3817 noise realisation.
@@ Appendix I KiDS-SBI posteriors and parameter constraints
To expand on the results of the KiDS-SBI analysis of the KiDS-1000 gold sample presented in Fig. 18 and Table 20, we present the full posterior learnt in KiDS-SBI to fit the data in Fig. 29 in Fig. 30 for both the anisotropic systematics and standard isotropic systematics models. Table 22 also provides the parameter best estimates for all cosmological and astrophysical parameters varied in the forward models. As shown in Fig. 30, both models leave $A_{\mathrm{bary}}$, $n_{\mathrm{s}}$ and $h_{0}$ unconstrained, while constraining $A_{\mathrm{IA}}$. The posteriors of both models strongly prefer values of $A_{\mathrm{IA}}$ greater than zero and less than 1.5. Despite the wide range of the $A_{\mathrm{IA}}$ priors, the posterior stays within a physical regime for the parameter which indicates a low chance of residual systematics strongly biasing the parameter constraints.
Figure 30: Posterior contours of the main constrained cosmological parameters from the KiDS-SBI analysis of the KiDS-1000 cosmic shear data assuming the anisotropic systematics model (in blue) compared against posterior contours from other analyses. In orange, we show the posterior from the same data while assuming the standard isotropic systematics model. In black and pink, we show the posteriors from the equivalent analyses done assuming a Gaussian likelihood for a given model. For clarity purposes, the $\bm{\delta}_{z}$ parameters are marginalised out, while $\omega_{\mathrm{b}}$ and $\omega_{\mathrm{c}}$ are folded into $\Omega_{\mathrm{m}}$. The green contour shows the posterior from the cosmic microwave background constraints from the TT,TE,EE+lowE modes (Planck Collaboration et al. 2020). We note that the Planck TT,TE,EE+lowE contours do not have any marginals in $A_{\mathrm{IA}}$ or $A_{\mathrm{bary}}$ as the CMB is not sensitive to the IAs of galaxies or baryonic feedback.